40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 7 19:42:01 2018, gautam, Update, General, IFO checkup and DRMI locking prep 
    Reply  Wed Nov 7 23:07:11 2018, gautam, Update, LSC, DRMI locking recovered Screenshot_from_2018-11-07_23-05-58.pngSRCL2MICH_crosscpl.pdfPRCangularCoh_rot.pdf
       Reply  Sat Nov 10 17:32:33 2018, gautam, Update, LSC, Nulling MICH->PRCL and MICH->SRCL sensMat_2018-11-10.pdf
          Reply  Tue Nov 13 18:09:24 2018, gautam, Update, LSC, Investigation of SRCL-->MICH coupling TF_sweptSineMeas.pdfdigitalDemod.pdfdigitalDemod_hist.pdfDRMI_LSCspectrogram.pdf
             Reply  Fri Nov 16 00:47:43 2018, gautam, Update, LSC, More DRMI characterization PRCff.pdfDRMI_withPRCff.png
Message ID: 14284     Entry time: Wed Nov 7 19:42:01 2018     Reply to this: 14285
Author: gautam 
Type: Update 
Category: General 
Subject: IFO checkup and DRMI locking prep 

Earlier today, I rebooted a few unresponsive VME crates (susaux, auxey).

The IMC has been unhappy for a couple of days - the glitches in the MC suspensions are more frequent. I reset the dark offsets, minimized MCREFL by hand, and then re-centered the beam on the MC2 Trans QPD. In this config, the IMC has been relatively stable today, although judging by the control room StripTool WFS control signal traces, the suspension glitches are still happening. Since we have to fix the attenuator issue anyways soon, we can do a touch-up on IMC WFS.

I removed the DC PD used for loss measurements. I found that the AS beam path was disturbed - there is a need to change the alignment, this just makes it more work to get back to IFO locking as I have to check alignment onto the AS55 and AS110 PDs.

Single arm locking worked with minimal effort - although the X arm dither alignment doesn't do the intended job of maximizing the transmission. Needs a checkup.

PRMI locking (carrier resonant) was also pretty easy. Stability of the lock is good, locks hold for ~20 minutes at a time and only broke because I was mucking around. However, when the carrier is resonant, I notice a smeared scatter pattern on the ITMX camera that I don't remember from before. I wonder if the FF idea can be tested in the simpler PRMI config.

After recovering these two simpler IFO configurations, I improved the cavity alignment by hand and with the ASS servos that work. Then I re-centered all the Oplev beams onto their respective QPDs and saved the alignment offsets. I briefly attemped DRMI locking, but had little success, I'm going to try a little later in the evening, so I'm leaving the IFO with the DRMI flashing about, LSC mode off.

ELOG V3.1.3-